OpenStudy (anonymous):

evaluate the indefinite integral.(4x-9)^-3(dx)

OpenStudy (luigi0210):

This? \[\LARGE \int\limits (4x-9)^{-3}~~dx\]

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

yes that's how it is wrote

OpenStudy (luigi0210):

Alrighty, give me a sec

OpenStudy (luigi0210):

We could use: \[\LARGE \int\limits x^n~dx=\frac{x^{n+1}}{n+1}+C \] right?

OpenStudy (anonymous):

yea not sure how

OpenStudy (zale101):

@satellite73

OpenStudy (dan815):